please i am dying help me out and solve this question with full explanation
14 is added 2/3 of a number. the result is 1 1/4 times the original number. find the number

Answer by rothauserc(4718)

let x be the number
:
2x/3 +14 = 5x/4
:
least common multiple is 12
:
multiply both sides of = by 12
:
8x +12 * 14 = 15x
:
8x +168 = 15x
:
subtract 8x from both sides of =
:
7x = 168
:
x = 168/7 = 24
:
****************
the number is 24
****************
:
check the answer
:
2(24)/3 +14 = 5(24)/4
:
16 +14 = 30
:
30 = 30
:
answer checks
: